# Optimal Best-Arm Identification under Fixed Confidence with Multiple Optima

### v1

@@ V-B Chernoff’s Stopping Rule
$Z_{a;b_{1},b_{2},\cdots,b_{M}}(t)$ $=N_{a}(t)d(\hat{\mu}_{a}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))+\sum_{i=1}^{M}N_{b_{i}}(t)d(\hat{\mu}_{b_{i}}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t)).$ (66)
We use the following stopping rule
$\tau_{\delta}$ $=\inf\bigg{\{}t\in\mathbb{N}:Z(t):=\max_{b_{1},b_{2},\cdots,b_{M}}\min_{a\in\mathcal{A}\setminus\{b_{1},b_{2},\cdots,b_{M}\}}Z_{a;b_{1},b_{2},\cdots,b_{M}}(t)>\beta(t,\delta)\bigg{\}},$ (67)
@@
  • Let $(\hat{b}_{1},\hat{b}_{2},\cdots,\hat{b}_{M})=\text{argmax}_{b_{1},b_{2},\cdots,b_{M}}\min_{a\in\mathcal{A}\setminus\{b_{1},b_{2},\cdots,b_{M}\}}Z_{a;b_{1},b_{2},\cdots,b_{M}}(\tau_{\delta})$.
@@ VI Choosing the threshold in the Stopping Rule
For simplicity, we only provide a proof for Bernoulli bandits (see (Garivier2016OptimalBA, )). The extension to other exponential bandits can be done by using the same methods as (Chambaz2005AMA, ).
@@
$\mathbb{P}_{\mu}\big{(}\tau_{\delta}<\infty,\hat{a}_{\tau_{\delta}}\notin\{1,2,\cdots,M\}\big{)}$
$\qquad\leq\mathbb{P}_{\mu}\Big{(}\exists\{b_{1},b_{2},\cdots,b_{M}\}\neq\{1,2,\cdots,M\},$
$\qquad\qquad\exists t\in\mathbb{N}:\min_{a\in\mathcal{A}\setminus\{b_{1},b_{2},\cdots,b_{M}\}}Z_{a;b_{1},b_{2},\cdots,b_{M}}(t)>\beta(t,\delta)\Big{)}.$ (68)
Now, for each tuple $(b_{1},b_{2},\cdots,b_{M})$ such that $\{b_{1},b_{2},\cdots,b_{M}\}\neq\{1,2,\cdots,M\}$, there exists
$a^{*}_{b_{1},b_{2},\cdots,b_{M}}\in\{1,2,\cdots,M\}\quad\text{such that}\quad a^{*}_{b_{1},b_{2},\cdots,b_{M}}\in\mathcal{A}\setminus\{b_{1},b_{2},\cdots,b_{M}\}.$
Then, we have
@@
$\mathbb{P}_{\mu}\big{(}\tau_{\delta}<\infty,\hat{a}_{\tau_{\delta}}\notin\{1,2,\cdots,M\}\big{)}$
@@
$1\leq e^{-\beta(t,\delta)}\frac{\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{2}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}{\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{2}}\wedge\cdots\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}.$ (70)
@@
$\leq\sum_{t=1}^{\infty}e^{-\beta(t,\delta)}\mathbb{E}_{\mu}\Bigg{[}\mathbbm{1}_{(T_{a;b_{1},b_{2},\cdots,b_{M}}=t)}\frac{\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}{\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}\Bigg{]}$ (72)
$\leq\sum_{t=1}^{\infty}e^{-\beta(t,\delta)}\mathbb{E}_{\mu}\Bigg{[}\mathbbm{1}_{(T_{a;b_{1},b_{2},\cdots,b_{M}}=t)}\frac{\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}{p_{\mu_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}\Bigg{]}.$ (73)
@@
×$\quad\times\underbrace{\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})\prod_{i\in\mathcal{A}\setminus\{a,b_{1},\cdots,b_{M}\}}p_{\mu_{i}}(\underline{x}_{t}^{i})}_{\text{joint density under alternative parameters}}dx_{1}dx_{2}\cdots dx_{t}.$ (74)
Now, by using (Garivier2016OptimalBA, , Lemma 11), we have
@@
$\qquad\leq 2\sqrt{n_{a}}\prod_{i=1}^{M}(2\sqrt{n_{b_{i}}})\,\text{kt}(\underline{x}_{t}^{a})\prod_{i=1}^{M}\text{kt}(\underline{x}_{t}^{b_{i}})$
$\qquad\leq 2^{M+1}\left(\frac{t}{M+1}\right)^{\frac{M+1}{2}}\text{kt}(\underline{x}_{t}^{a})\prod_{i=1}^{M}\text{kt}(\underline{x}_{t}^{b_{i}}),$ (75)
where $\text{kt}(\cdot)$ is the Krichevsky–Trofimov distribution (see (Krichevsky1981, )).
Therefore,
@@
$\leq\sum_{t=1}^{\infty}2^{M+1}\left(\frac{t}{M+1}\right)^{\frac{M+1}{2}}e^{-\beta(t,\delta)}\int_{\{0,1\}^{t}}\mathbbm{1}_{(T_{a;b_{1},b_{2},\cdots,b_{M}}=t)}(x_{1},\cdots,x_{t})$
×$\quad\times\text{kt}(\underline{x}_{t}^{a})\prod_{i=1}^{M}\text{kt}(\underline{x}_{t}^{b_{i}})\prod_{i\in\mathcal{A}\setminus\{a,b_{1},\cdots,b_{M}\}}p_{\mu_{i}}(\underline{x}_{t}^{i})\,dx_{1}\cdots dx_{t}.$ (76)
@@
$\mathbb{P}_{\mu}\big{(}T_{a;b_{1},b_{2},\cdots,b_{M}}<\infty\big{)}$ $\leq\sum_{t=1}^{\infty}2^{M+1}\left(\frac{t}{M+1}\right)^{\frac{M+1}{2}}e^{-\beta(t,\delta)}\tilde{\mathbb{P}}\big{(}T_{a;b_{1},b_{2},\cdots,b_{M}}=t\big{)}$
$\leq\frac{\delta}{K^{M}}\tilde{\mathbb{P}}\big{(}T_{a;b_{1},b_{2},\cdots,b_{M}}<\infty\big{)}\leq\frac{\delta}{K^{M}}.$ (77)

### v2

@@ V-B Chernoff’s Stopping Rule
    $Z_{a;b_{1},b_{2},\cdots,b_{M}}(t)=N_{a}(t)d(\hat{\mu}_{a}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))$
    $\quad+\sum_{i=1}^{M}N_{b_{i}}(t)d(\hat{\mu}_{b_{i}}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t)).$ (70)
  • Case 2: $\mathcal{V}\neq\emptyset$. Then, for all $i\in\mathcal{U}$, we have $\mu^{\prime}_{b_{i}}=\mu^{\prime}_{a}$ by (64). Then, from (63) we have
    $\sum_{i\in\mathcal{U}}N_{b_{i}}(t)\big(\hat{\mu}_{b_{i}}(t)-\mu^{\prime}_{b_{i}}\big)\bigg[\frac{1}{\ddot{A}(\dot{A}^{-1}(\mu^{\prime}_{b_{i}}))}\bigg]-\beta_{i}=0.$ (71)
    Hence, from (62) and (71) we obtain
    $N_{a}(t)\big(\hat{\mu}_{a}(t)-\mu^{\prime}_{a}\big)\bigg[\frac{1}{\ddot{A}(\dot{A}^{-1}(\mu^{\prime}_{a}))}\bigg]$
    $\quad+\sum_{i\in\mathcal{U}}N_{b_{i}}(t)\big(\hat{\mu}_{b_{i}}(t)-\mu^{\prime}_{a}\big)\bigg[\frac{1}{\ddot{A}(\dot{A}^{-1}(\mu^{\prime}_{a}))}\bigg]=0.$ (72)
    From (72) we obtain
    $\mu^{\prime}_{a}=\frac{N_{a}(t)\hat{\mu}_{a}(t)+\sum_{i\in\mathcal{U}}N_{b_{i}}(t)\hat{\mu}_{b_{i}}(t)}{N_{a}(t)+\sum_{i\in\mathcal{U}}N_{b_{i}}(t)}:=\bar{\mu}_{a}(t).$ (73)
    Now, from (63) we obtain
    $\mu^{\prime}_{b_{i}}=\hat{\mu}_{b_{i}}(t),\qquad\forall i\in\mathcal{B}.$ (74)
    Then, by (65) and (74) we have
    $\bar{\mu}_{a}(t)\geq\max_{i\in\mathcal{V}}\mu^{\prime}_{b_{i}}=\max_{i\in\mathcal{V}}\hat{\mu}_{b_{i}}(t)\geq\frac{\sum_{i\in\mathcal{V}}N_{b_{i}}(t)\hat{\mu}_{b_{i}}(t)}{\sum_{i\in\mathcal{V}}N_{b_{i}}(t)}.$ (75)
    Now, using the fact that for any $p,q,r,s>0$ if $\frac{p}{q}\leq\frac{r}{s}$ we have $\frac{p}{q}\leq\frac{p+r}{q+s}\leq\frac{r}{s}$, from (73) and (75) we obtain
    $\frac{\sum_{i\in\mathcal{V}}N_{b_{i}}(t)\hat{\mu}_{b_{i}}(t)}{\sum_{i\in\mathcal{V}}N_{b_{i}}(t)}\leq\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t)\leq\bar{\mu}_{a}(t).$ (76)
    Observe that
    $\frac{\sum_{i\in\mathcal{V}}N_{b_{i}}(t)\hat{\mu}_{b_{i}}(t)}{\sum_{i\in\mathcal{V}}N_{b_{i}}(t)}\geq\min_{i\in\mathcal{V}}\hat{\mu}_{b_{i}}(t)\geq\min_{i\in[M]}\hat{\mu}_{b_{i}}(t).$ (77)
    On the other hand, from (63) and $\beta_{i}\geq 0\kern 5.0pt\forall i\in[M]$, we have
    $\bar{\mu}_{a}(t)=\mu^{\prime}_{a}\leq\min_{i\in\mathcal{U}}\hat{\mu}_{b_{i}}(t).$ (78)
    It follows from (76), (77), and (78) that
    $\min_{i\in[M]}\hat{\mu}_{b_{i}}(t)\leq\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t)\leq\bar{\mu}_{a}(t)$
    $\quad\leq\min_{i\in\mathcal{U}}\hat{\mu}_{b_{i}}(t)\leq\max_{i\in[M]}\hat{\mu}_{b_{i}}(t).$ (79)
    Therefore, we obtain
    $\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t)\leq\bar{\mu}_{a}(t)\leq\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t)+\zeta(t),$ (80)
    where
    $\zeta(t):=\max_{i\in[M]}\hat{\mu}_{b_{i}}(t)-\min_{i\in[M]}\hat{\mu}_{b_{i}}(t).$ (81)
Hence, we have
$\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{2}}\wedge\cdots\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}\bigg[l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})\bigg]$
$\quad=l_{a}(\bar{\mu}_{a}(t))+\sum_{i\in\mathcal{U}}l_{b_{i}}(\bar{\mu}_{a}(t))+\sum_{i\in\mathcal{V}}l_{b_{i}}(\hat{\mu}_{b_{i}}(t)).$ (82)
Under the condition $\hat{\mu}_{a}(t)\leq\min_{i\in[M]}\hat{\mu}_{b_{i}}(t)$, it holds that
$\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t)\geq\hat{\mu}_{a}(t).$ (83)
Hence, by using the fact that $l_{x}(\mu^{\prime}_{x})$ is non-increasing if $\mu^{\prime}_{x}\in[\hat{\mu}_{x}(t),\infty)$ and non-decreasing if $\mu^{\prime}_{x}\in(-\infty,\hat{\mu}_{x}(t)])$, from (82) and (80) we obtain
$\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}\bigg[l_{a}(\mu^{\prime}_{a})+\sum_{i=1}^{M}l_{b_{i}}(\mu^{\prime}_{b_{i}})\bigg]$
$\quad\geq l_{a}(\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t)+\zeta(t))$
$\qquad+\sum_{i\in\mathcal{U}}l_{b_{i}}(\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))+\sum_{i\in\mathcal{V}}l_{b_{i}}(\hat{\mu}_{b_{i}}(t)).$ (84)
$Z_{a;b_{1},b_{2},\cdots,b_{M}}(t)$
$\quad\leq N_{a}(t)d(\hat{\mu}_{a}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t)+\zeta(t))$
$\qquad+\sum_{i\in\mathcal{U}}N_{b_{i}}(t)d(\hat{\mu}_{b_{i}}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))$ (85)
$\quad=N_{a}(t)d(\hat{\mu}_{a}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))$
$\qquad+\sum_{i=1}^{M}N_{b_{i}}(t)d(\hat{\mu}_{b_{i}}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))+R(t),$ (86)
where
$R(t)=N_{a}(t)\big[d(\hat{\mu}_{a}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t)+\zeta(t))$
$\quad-d(\hat{\mu}_{a}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))\big]$
$\qquad-\sum_{i\in\mathcal{V}}N_{b_{i}}(t)d(\hat{\mu}_{b_{i}}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))$ (87)
$\quad\leq t\bigg(\frac{N_{a}(t)}{t}\big[d(\hat{\mu}_{a}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t)+\zeta(t))$
$\qquad-d(\hat{\mu}_{a}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))\big]$
$\qquad-\min_{i\in[M]}\frac{N_{b_{i}}(t)}{t}d(\hat{\mu}_{b_{i}}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))\bigg)$ (88)
$\quad\leq 0$ (89)
under the condition that
$\frac{N_{a}(t)}{t}\big[d(\hat{\mu}_{a}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t)+\zeta(t))$
$\quad-d(\hat{\mu}_{a}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))\big]$
$\qquad-\min_{i\in[M]}\frac{N_{b_{i}}(t)}{t}d(\hat{\mu}_{b_{i}}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))\leq 0.$ (90)
From (70) and (89), we finally obtain
$Z_{a;b_{1},b_{2},\cdots,b_{M}}(t)$
$\quad=N_{a}(t)d(\hat{\mu}_{a}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))$
$\qquad+\sum_{i=1}^{M}N_{b_{i}}(t)d(\hat{\mu}_{b_{i}}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t)),$ (91)
under the condition that $(\hat{\mu}_{a}(t)\leq\hat{\mu}_{b_{1}}(t))\wedge(\hat{\mu}_{a}(t)\leq\hat{\mu}_{b_{2}}(t))\wedge\cdots\wedge(\hat{\mu}_{a}(t)\leq\hat{\mu}_{b_{M}}(t))$ and
$S_{a;b_{1},b_{2},\cdots,b_{M}}(t)$
$\quad:=\frac{N_{a}(t)}{t}\big[d(\hat{\mu}_{a}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t)+\zeta(t))$
$\qquad-d(\hat{\mu}_{a}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))\big]$
$\qquad-\min_{i\in[M]}\frac{N_{b_{i}}(t)}{t}d(\hat{\mu}_{b_{i}}(t),\hat{\mu}_{a;b_{1},b_{2},\cdots,b_{M}}(t))\leq 0.$ (92)
Next, we use the following stopping rule
$\tau$ $=\inf\bigg\{t\in\mathbb{N}:Z(t):=\max_{b_{1},b_{2},\cdots,b_{M}}\min_{a\in\mathcal{A}\setminus\{b_{1},b_{2},\cdots,b_{M}\}}$
$\qquad Z_{a;b_{1},b_{2},\cdots,b_{M}}(t)>\beta(t,\delta)\bigg\},$ (93)
@@
  • Let
    $(\hat{b}_{1},\hat{b}_{2},\cdots,\hat{b}_{M})=\text{argmax}_{b_{1},b_{2},\cdots,b_{M}}\min_{a\in\mathcal{A}\setminus\{b_{1},b_{2},\cdots,b_{M}\}}$
    $\quad Z_{a;b_{1},b_{2},\cdots,b_{M}}(\tau).$ (94)
@@ VI Choosing the threshold in the Stopping Rule
For simplicity, we only provide a proof for Bernoulli bandits (see [3]). The extension to other exponential bandits can be done by using the same methods as [12].
@@
$\mathbb{P}_{\mu}\big(\tau<\infty,\hat{a}_{\tau}\notin[M]\big)$
$\quad\leq\mathbb{P}_{\mu}\Big(\exists\{b_{1},b_{2},\cdots,b_{M}\}\neq[M],$
$\quad\quad\exists t\in\mathbb{N}:\min_{a\in\mathcal{A}\setminus\{b_{1},b_{2},\cdots,b_{M}\}}Z_{a;b_{1},b_{2},\cdots,b_{M}}(t)>\beta(t,\delta)\Big).$ (95)
Now, for each tuple $(b_{1},b_{2},\cdots,b_{M})$ such that $\{b_{1},b_{2},\cdots,b_{M}\}\neq[M]$, there exists $a^{*}_{b_{1},b_{2},\cdots,b_{M}}\in[M]\quad\text{such that}\quad a^{*}_{b_{1},b_{2},\cdots,b_{M}}\in\mathcal{A}\setminus\{b_{1},b_{2},\cdots,b_{M}\}.$ Then, we have
@@
$\mathbb{P}_{\mu}\big(\tau<\infty,\hat{a}_{\tau}\notin[M]\big)$
@@
$1\leq e^{-\beta(t,\delta)}$
$\times\frac{\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}{\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}\Bigg].$ (97)
@@
$\leq\sum_{t=1}^{\infty}e^{-\beta(t,\delta)}\mathbb{E}_{\mu}\Bigg[\mathbbm{1}_{(T_{a;b_{1},b_{2},\cdots,b_{M}}=t)}$
$\quad\frac{\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}{\max_{\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\geq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}\Bigg]$ (99)
$\leq\sum_{t=1}^{\infty}e^{-\beta(t,\delta)}\mathbb{E}_{\mu}\Bigg[\mathbbm{1}_{(T_{a;b_{1},b_{2},\cdots,b_{M}}=t)}$
$\quad\frac{\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}{p_{\mu_{a}}(\underline{X}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu_{b_{i}}}(\underline{X}_{N_{b_{i}}(t)}^{b_{i}})}\Bigg].$ (100)
@@
$\quad\times\max_{\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{1}}\wedge\cdots\wedge\mu^{\prime}_{a}\leq\mu^{\prime}_{b_{M}}}p_{\mu^{\prime}_{a}}(\underline{x}_{N_{a}(t)}^{a})\prod_{i=1}^{M}p_{\mu^{\prime}_{b_{i}}}(\underline{x}_{N_{b_{i}}(t)}^{b_{i}})$
$\qquad\prod_{i\in\mathcal{A}\setminus\{a,b_{1},\cdots,b_{M}\}}p_{\mu_{i}}(\underline{x}_{N_{i}(t)}^{i})dx_{1}dx_{2}\cdots dx_{t}.$ (101)
Now, by using [3, Lemma 11], we have
@@
$\quad\leq 2\sqrt{N_{a}(t)}\prod_{i=1}^{M}(2\sqrt{N_{b_{i}}(t)})\,\text{kt}(\underline{x}_{N_{a}(t)}^{a})\prod_{i=1}^{M}\text{kt}(\underline{x}_{N_{b_{i}}(t)}^{b_{i}})$
$\quad\leq 2^{M+1}\left(\frac{t}{M+1}\right)^{\frac{M+1}{2}}\text{kt}(\underline{x}_{N_{a}(t)}^{a})\prod_{i=1}^{M}\text{kt}(\underline{x}_{N_{b_{i}}(t)}^{b_{i}}),$ (102)
where $\text{kt}(\cdot)$ is the Krichevsky–Trofimov distribution (see [13]). Here, we use the fact that $\nu_{1}\nu_{2}\cdots\nu_{M+1}\leq(\frac{\nu_{1}+\nu_{2}+\cdots+\nu_{M+1}}{M+1})^{M+1}$ for all $\nu_{1},\nu_{2},\cdots,\nu_{M+1}\geq 0$ in (102).
Therefore, we obtain
@@
$\quad\leq\sum_{t=1}^{\infty}2^{M+1}\left(\frac{t}{M+1}\right)^{\frac{M+1}{2}}e^{-\beta(t,\delta)}$
$\qquad\times\int_{\{0,1\}^{t}}\mathbbm{1}_{(T_{a;b_{1},b_{2},\cdots,b_{M}}=t)}(x_{1},\cdots,x_{t})$
$\qquad\times\text{kt}(\underline{x}_{N_{a}(t)}^{a})\prod_{i=1}^{M}\text{kt}(\underline{x}_{N_{b_{i}}(t)}^{b_{i}})$
$\qquad\times\prod_{i\in\mathcal{A}\setminus\{a,b_{1},\cdots,b_{M}\}}p_{\mu_{i}}(\underline{x}_{N_{i}(t)}^{i})\,dx_{1}\cdots dx_{t}.$ (103)
@@
$\mathbb{P}_{\mu}\big(T_{a;b_{1},b_{2},\cdots,b_{M}}<\infty\big)$
$\leq\sum_{t=1}^{\infty}2^{M+1}\left(\frac{t}{M+1}\right)^{\frac{M+1}{2}}e^{-\beta(t,\delta)}\tilde{\mathbb{P}}\big(T_{a;b_{1},b_{2},\cdots,b_{M}}=t\big)$
$\leq\frac{\delta}{K^{M}}\tilde{\mathbb{P}}\big(T_{a;b_{1},b_{2},\cdots,b_{M}}<\infty\big)$
$\leq\frac{\delta}{K^{M}}.$ (104)